# Integrating Low-Altitude SAR Imaging into UAV Data Backhaul

### v1

@@ Abstract
Synthetic aperture radar (SAR) deployed on unmanned aerial vehicles (UAVs) is expected to provide burgeoning imaging services for low-altitude wireless networks (LAWNs), thereby enabling large-scale environmental sensing and timely situational awareness. Conventional SAR systems typically leverages a deterministic radar waveform, while it conflicts with the integrated sensing and communications (ISAC) paradigm by discarding signaling randomness, in whole or in part. In fact, this approach reduces to the uplink pilot sensing in 5G New Radio (NR) with sounding reference signals (SRS), underutilizing data symbols. To explore the potential of data-aided imaging, we develop a low-altitude SAR imaging framework that sufficiently leverages data symbols carried by the native orthogonal frequency division multiplexing (OFDM) communication waveform. The randomness of modulated data in the temporal-frequency (TF) domain, introduced by non-constant modulus constellations such as quadrature amplitude modulation (QAM), may however severely degrade the imaging quality. To mitigate this effect, we incorporate several TF-domain filtering schemes within a range-Doppler (RD) imaging framework and evaluate their impact. We further propose using the normalized mean square error (NMSE) of a reference point target’s profile as an imaging performance metric. Simulation results with 5G NR parameters demonstrate that data-aided imaging substantially outperforms pilot-only counterpart, accordingly validating the effectiveness of the proposed OFDM-SAR imaging approach in LAWNs.
@@ I Introduction
With the rapid advances of the low-altitude economy, unmanned aerial vehicles (UAVs) are increasingly deployed as key platforms for civilian and industrial missions such as urban inspection, logistics, infrastructure monitoring, and environmental observation. Benefiting from their mobility, low cost, and flexible deployment in low-altitude wireless networks (LAWNs), UAVs provide an efficient means for large-scale data collection and real-time situational awareness [wu2025low, wu2025toward, song2025overview]. In particular, when equipped with synthetic aperture radar (SAR) sensors [cumming2005digital, richards2005fundamentals, wang2013mimo, zhang2024optical], UAVs are capable of providing high-resolution and all-weather imaging that is robust to illumination and atmospheric variations. Therefore, UAV-SAR in LAWNs enables accurate environmental perception and mapping, which are essential for urban management, intelligent transportation, and public security.
@@
However, in many time-critical missions, UAV-SAR imaging must satisfy harsh real-time requirements. For instance, in urban traffic monitoring and moving target tracking, continuous imaging and low-latency data backhaul are crucial for adaptive control and rapid decision-making [du2025toward]. This imposes dual challenges on UAV-SAR systems: the need for precise, high-resolution imaging, and the requirement for high-throughput, low-latency data transmission. To this end, communication and sensing modules on current UAV platforms are typically implemented separately. The sensing module integrates various sensors such as radar, LiDAR, and cameras, which occupy significant space and weight while introducing considerable power consumption, thereby limiting the UAV’s endurance. To address this issue, integrating radar sensing and communications into a unified module is a promising solution. Current literature exploit a unified integrated sensing and communications (ISAC) signaling strategies, mainly focusing on sensing-centric [wang2019first, chatzitheodoridi2022cooperative] and joint designs [yang2022waveform]. In contrast, communication-centric designs [sturm2011waveform, kumari2017ieee, cui2021integrating], which has received comparatively less attention in SAR imaging, may offer a potentially more practical and cost-effective solution. By reusing the existing communication hardware for imaging purposes, this approach not only improves spectrum efficiency but also helps mitigate mutual interference between the imaging and communication functions. In this spirit, joint orthogonal frequency-division multiplexing (OFDM) SAR imaging and data transmission has emerged as a promising approach. By leveraging the OFDM waveform, the system can simultaneously achieve high-quality SAR imaging and communications, thus formulating a new ISAC paradigm [liu2022integrated] in LAWNs.
@@
Standardized OFDM has been widely employed as the default waveform in cellular communications[hwang2008ofdm] and WiFi [niu2022rethinking], due to its high spectral efficiency and robustness against multipath fading. In addition, the OFDM waveform has been increasingly employed in radar systems for target detection, estimation and tracking tasks [berger2010signal, shi2017power, bicua2016generalized, sen2010adaptive, du2020distributed], owing to its flexible spectrum design and thumbtack-shaped ambiguity function. In particular, its linear temporal-frequency (TF) structure allows straightforward delay-Doppler information extraction using 2D-fast Fourier transform (FFT), while the embedding of cyclic prefix (CP) provides robustness against frequency-selective fading [du2025prob]. In addition to the sensing functionalities above, the radar imaging capability of exploiting OFDM signals, has also been frequently reported in the last decade [garmatyuk2003ultra, garmatyuk2011adaptive, zhang2014ofdm, zhang2014irci, yu2018irci, wang2019first]. The motivation also benefits from the exploitation of multicarrier structure and the CP, which can readily synthesize a large bandwidth signaling, and mitigate the inter-range cell interference (IRCI) induced by the inter-symbol interference (ISI).
@@
In spite of this, aforementioned works replace the random data of OFDM communications signals with time-repeated deterministic/pseudo noise-like weighting coefficients [garmatyuk2003ultra, garmatyuk2011adaptive, zhang2014ofdm, zhang2014irci, yu2018irci], or multiplex sensing and communications in different resource blocks [wang2019first]. As such, the data transmission capability is significantly sacrificed. Notably, the above sensing signaling may be treated as pilots in the context of 5G New Radio (NR) frame structure, such as the sounding reference signals (SRS) [dahlman20205g] in the uplink channel. However, SRS occupies only a small fraction of TF resources, resulting in limited bandwidth, pulse duration, and energy accumulation, which in turn degrades imaging resolution and accuracy. Moreover, the low pulse repetition frequency (PRF) of pilot transmissions may lead to Doppler aliasing due to reduced azimuth sampling density, thereby impairing the azimuth focusing performance. In addition, SRS has a discontinuous comb structure across subcarriers [etsi38138], which leads to periodic peaks in the time domain and thus deteriorate the imaging performance. On top of these factors, most downlink signals from ground base stations to UAVs merely convey control instructions, while the UAV must upload large volumes of imaging data, leading to a much higher uplink backhaul throughput. To comprehensively balance imaging and communication qualities, together with computational complexities, the communication-centric signaling design [sturm2011waveform, kumari2017ieee, cui2021integrating] maintaining compatibility with the existing 5G NR protocol [dahlman20205g] is thus envisioned as a more economically viable paradigm with data-aided symbols for imaging.
@@
Nevertheless, data-aided imaging triggers a new challenge: the randomness must be effectively mitigated to restore the orthogonality between the range and azimuth steering vectors through proper TF-domain filtering schemes [du2024reshaping], thereby ensuring compatibility with conventional 2D Fourier-based imaging algorithms such as range-Doppler (RD) and Omega-$K$ methods [cumming2005digital]. To be specific, data symbols are typically drawn from quadrature amplitude modulation (QAM) constellations, which are differ both in time and frequency dimensions [du2024reshaping]. Consequently, the range and azimuth profiles may be simultaneously jeopardized by the signaling randomness, leading to compromised range-azimuth compression performance. To mitigate this effect, data-dependent filtering can be performed in the TF domain, such as matched filtering [kay1993fundamentals], element wise division [sturm2011waveform, zhang2014ofdm] or linear minimum mean squared error (LMMSE) filtering [du2025prob, keskin2024fundamental]. However, a quantitative characterization of imaging performance under data-aided scheme remains unclear, particularly in terms of metrics that can capture the impact of signaling randomness and filtering design.
@@
  • Unlike conventional OFDM-SAR designs inspired by 5G NR, which effectively exploit only pilot signals (e.g., SRS) for sensing, we establish a data-aided imaging model that explicitly incorporates data symbols into the imaging process. Since the inherent randomness of communication data can noticeably distort the echo statistics and degrade imaging quality, the proposed TF-domain filters are designed to suppress such randomness, thereby regularizing the echo structure and enabling the resulting echoes compatible with the standard RD algorithm.
  • Furthermore, we propose a quantitative framework to evaluate the quality of data-aided SAR imaging, based on the normalized mean square error (NMSE) of the range-azimuth profiles with respect to a reference point target, such as a corner reflector commonly used in practical SAR measurements. Compared with relying solely on conventional metrics such as the integrated sidelobe ratio (ISLR) [stoica2009new], noise-equivalent sigma zero (NESZ) [patyuchenko2009design], and peak energy loss (PEL) [du2025prob], the NMSE provides a more comprehensive evaluation of imaging quality.
@@
Notation: Throughout the paper, $\bm{A}$, $\bm{a}$, and $a$ denote a matrix, vector, and scalar, respectively; $|a|$, $|\bm{A}|$ and $\frac{\bm{B}}{\bm{A}}$ represent the modulus of $a$, the element-wise modulus of $\bm{A}$ and the element-wise division of $\bm{B}$ by $\bm{A}$, respectively; $\mathbb{E}(\cdot)$, $tr(\cdot)$, $\|\cdot\|_{F}$, $(\cdot)^{T}$, $(\cdot)^{*}$, $(\cdot)^{H}$, $\odot$ and $\bm{I}$ denote the expectation, trace, Frobenius norm, transpose, conjugate, Hermitian, Hadamard (element-wise) product and identity matrix, respectively; A complex Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$ is represented as $\mathcal{CN}(\mu,\sigma^{2})$; Finally, $\mathrm{sinc}(x)=\frac{\sin(\pi x)}{\pi x}$.
Figure 1: Joint OFDM-SAR imaging and data backhaul.
@@ II-A SAR Geometry
As depicted in Fig. 1, we consider a joint monostatic broadside stripmap UAV-SAR imaging embedded in continuous data backhaul scheme with ground cellular base stations in LAWNs. The sensor platform moves along the $y$-axis, with a height of $H_{p}$. To guarantee uninterrupted communications, the UAV’s coverage footprint must exceed the radius of a base station, ensuring at least one base station remains in line-of-sight of SAR beampattern at all times. This is particularly important as the sensor sweeps a wide swath along its flight path, resulting in varying azimuth positions relative to base stations. Consequently, a quantitatively feasibility analysis is required.
Let us commence by analyzing the azimuth beamwidth and the elevation beamwidth, expressed as [cumming2005digital]
@@
where $\lambda$ is the wavelength, $D_{\mathrm{al}}$ and $D_{\mathrm{el}}$ represent the azimuth and elevation antenna aperture, respectively. With a typical height $H_{p}$ from $300$ to $3000$ meters in LAWNs, the center of SAR elevation beamwidth defines the beam’s pointing elevation angle $\theta_{c}$. Accordingly, azimuth coverage length $L_{\mathrm{az}}$ and the elevation coverage length $L_{\mathrm{el}}$ can be approximated as
@@
As clarified above, in order to guarantee the real-time data backhaul, $L_{\mathrm{az}}$ and $L_{\mathrm{el}}$ must be larger than the cellular coverage. For example, we consider $3.5$ GHz carrier frequency in sub-6 GHz band, azimuth and elevation antenna apertures as $D_{\mathrm{az}}=D_{\mathrm{el}}=0.1$ m, a platform height $H_{p}=1000$ m, and a central elevation angle $\theta_{c}=\frac{\pi}{4}$. In this case, the corresponding ground coverage lengths are approximately $L_{\mathrm{az}}\approx 564$ m and $L_{\mathrm{el}}\approx 1899$ m, which obviously exceed the coverage of a 5G cellular cell.
Note that there is a trade-off between these parameters. Specifically, a higher altitude leads to a larger beam footprint, while it also implies a larger slant range, which requires a longer CP duration to avoid ISI and inter-carrier interference (ICI) [du2025prob, keskin2024fundamental, keskin2021mimo]. Therefore, the system parameters should be carefully devised by referring to 5G NR. Based on these parameter analysis, the practical feasibility of integrating OFDM-SAR imaging into UAV data backhaul in LAWNs is geometrically convinced.
@@ II-B Transmit ISAC Signal Model
In (5), $s_{n,m}$ denotes the transmitted frequency-domain data, drawn from a finite alphabet such as 256-QAM constellation. In addition, $\Delta f=B_{r}/N=1/T_{p}$ represents the subcarrier interval in the frequency domain, and $\mathrm{rect}(t)$ represents the rectangle window, equal to $1$ for $0\leq t\leq 1$, and zero otherwise [^1] Note that the practical SAR has a sinc-like azimuth envelope [cumming2005digital, zhang2014ofdm]. For the theoretical convenience, we use an ideal rectangle envelope instead.. Additionally, $T_{\mathrm{sym}}=T_{p}+T_{cp}$, where $T_{cp}$ denotes the length of CP. In order to eliminate the ISI and the ICI of received sensing signals, $T_{cp}$ must be larger than round-trip delay of the farthest target/path. The total azimuth time is thus $T_{a}=MT_{\mathrm{sym}}$. Note that the power of a given codebook is normalized, i.e., $\mathbb{E}\{|s_{n,m}|^{2}\}=1$ for $\forall n,m$. This leads to the fact that the total average transmit energy of $M$ symbols is also normalized as $\mathbb{E}\{|x(t)|^{2}\}=\frac{1}{NM}\mathbb{E}\left\{\sum_{n,m}|s_{n,m}|^{2}\right\}=1$.

### v2

@@ Abstract
Synthetic aperture radar (SAR) on unmanned aerial vehicles (UAVs) enables high-resolution sensing in low-altitude wireless networks, while requiring reliable uplink data backhaul to ground base stations under dynamic channel conditions. Conventional orthogonal frequency division multiplexing (OFDM)-based SAR systems rely on pilot or deterministic signaling, which occupies only a small fraction of the available time-frequency (TF) resources and limits imaging performance. This paper develops a data-aided OFDM-SAR imaging framework that reuses uplink communication data symbols for sensing, thereby exploiting the dominant TF resources of the UAV backhaul link. However, the randomness of data symbols disrupts the coherent structure required for SAR imaging, especially in highly dynamic channels with strong TF coupling, leading to severe degradation in range-Doppler focusing. To address this issue, we establish a unified TF domain filtering framework to suppress data-induced randomness and recover an equivalent deterministic imaging channel. Within this framework, reciprocal, matched, and Wiener filtering are interpreted under a common formulation, enabling a systematic characterization of their impact on imaging performance. A normalized mean square error (NMSE) metric of a reference point target’s profile is further adopted to quantify the joint effects of randomness-induced distortion and noise amplification. Simulation results based on 5G NR parameters show that the proposed data-aided scheme significantly outperforms pilot-only approaches by leveraging uplink data resources, demonstrating that effective TF-domain filtering is essential to ensure high-resolution imaging in dynamic UAV channels.
@@ I Introduction
With the rapid advances of the low-altitude economy, unmanned aerial vehicles (UAVs) are increasingly deployed as key platforms for civilian and industrial missions such as urban inspection, logistics, infrastructure monitoring, and environmental observation. Benefiting from their mobility, low cost, and flexible deployment in low-altitude wireless networks (LAWNs), UAVs provide an efficient means for large-scale data collection and real-time situational awareness [40, 41, 32, 25]. In particular, when equipped with synthetic aperture radar (SAR) sensors [5, 29, 37, 46], UAVs are capable of providing high-resolution and all-weather imaging that is robust to illumination and atmospheric variations. Therefore, UAV-SAR in LAWNs enables accurate environmental perception and mapping, which are essential for urban management, intelligent transportation, and public security.
@@
However, in many time-critical missions, UAV-SAR imaging must satisfy harsh real-time requirements. For instance, in urban traffic monitoring and moving target tracking, continuous imaging and low-latency data backhaul are crucial for adaptive control and rapid decision-making [7]. This imposes dual challenges on UAV-SAR systems: the need for precise, high-resolution imaging, and the requirement for high-throughput, low-latency data transmission. To this end, communication and sensing modules on current UAV platforms are typically implemented separately. The sensing module integrates various sensors such as radar, LiDAR, and cameras, which occupy significant space and weight while introducing considerable power consumption, thereby limiting the UAV’s endurance. To address this issue, integrating radar sensing and communications into a unified module is a promising solution. Current literature exploit a unified integrated sensing and communications (ISAC) signaling strategies, mainly focusing on sensing-centric [36, 3, 15] and joint designs [42]. In contrast, communication-centric designs [34, 20, 4, 35, 24], which has received comparatively less attention in SAR imaging, may offer a potentially more practical and cost-effective solution. By reusing the existing communication hardware for imaging purposes, this approach not only improves spectrum efficiency but also helps mitigate mutual interference between the imaging and communication functions. In this spirit, joint orthogonal frequency-division multiplexing (OFDM) SAR imaging and data transmission has emerged as a promising approach. By leveraging the OFDM waveform, the system can simultaneously achieve high-quality SAR imaging and communications, thus formulating a new ISAC paradigm [21] in LAWNs.
@@
Standardized OFDM has been widely employed as the default waveform in cellular communications[16] and WiFi [27], due to its high spectral efficiency and robustness against multipath fading. In addition, the OFDM waveform has been increasingly employed in radar systems for target detection, estimation and tracking tasks [1, 31, 2, 30, 10], owing to its flexible spectrum design and thumbtack-shaped ambiguity function. In particular, its linear temporal-frequency (TF) structure allows straightforward delay-Doppler information extraction using 2D-fast Fourier transform (FFT), while the embedding of cyclic prefix (CP) provides robustness against frequency-selective fading [9]. In addition to the sensing functionalities above, the radar imaging capability of exploiting OFDM signals, has also been frequently reported in the last decade [14, 13, 45, 44, 43, 36]. The motivation also benefits from the exploitation of multicarrier structure and the CP, which can readily synthesize a large bandwidth signaling, and mitigate the inter-range cell interference (IRCI) induced by the inter-symbol interference (ISI).
@@
In spite of this, aforementioned works replace the random data of OFDM communications signals with time-repeated deterministic/pseudo noise-like weighting coefficients [14, 13, 45, 44, 43], or multiplex sensing and communications in different resource blocks [36]. As such, the data transmission capability is significantly sacrificed. Notably, the above sensing signaling may be treated as pilots in the context of 5G New Radio (NR) frame structure, such as the sounding reference signals (SRS) [6] in the uplink channel. However, SRS occupies only a small fraction of TF resources, resulting in limited bandwidth, pulse duration, and energy accumulation, which in turn degrades imaging resolution and accuracy. Moreover, the low pulse repetition frequency (PRF) of pilot transmissions may lead to Doppler aliasing due to reduced azimuth sampling density, thereby impairing the azimuth focusing performance. In addition, SRS has a discontinuous comb structure across subcarriers [12], which leads to periodic peaks in the time domain and thus deteriorate the imaging performance. On top of these factors, most downlink signals from ground base stations to UAVs merely convey control instructions, while the UAV must upload large volumes of imaging data, leading to a much higher uplink backhaul throughput. To comprehensively balance imaging and communication qualities, together with computational complexities, the communication-centric signaling design [34, 20, 4] maintaining compatibility with the existing 5G NR protocol [6] is thus envisioned as a more economically viable paradigm with data symbols for imaging.
@@
Nevertheless, data-aided imaging triggers a new challenge: the randomness must be effectively mitigated to restore the orthogonality between the range and azimuth steering vectors through proper TF-domain filtering schemes [8], thereby ensuring compatibility with conventional 2D Fourier-based imaging algorithms such as range-Doppler (RD) and Omega-$K$ methods [5]. To be specific, data symbols are typically drawn from quadrature amplitude modulation (QAM) constellations, which are differ both in time and frequency dimensions [8]. Consequently, the range and azimuth profiles may be simultaneously jeopardized by the signaling randomness, leading to compromised range-azimuth compression performance. To mitigate this effect, data-dependent filtering can be performed in the TF domain, such as matched filtering [17], element wise division [34, 45] or linear minimum mean squared error (LMMSE) filtering [9, 18]. However, a quantitative characterization of imaging performance under data-aided scheme remains unclear, particularly in terms of metrics that can capture the impact of signaling randomness and filtering design.
@@
  • Unlike conventional OFDM-SAR designs inspired by 5G NR, which effectively exploit only pilot signals (e.g., SRS) for sensing, we establish a data-aided imaging model that explicitly incorporates data symbols into both range and azimuth dimensions. Since the inherent randomness of communication data can noticeably distort the echo statistics and degrade imaging quality, the proposed TF-domain filters are designed to suppress such randomness, thereby regularizing the echo structure and enabling it compatible with the standard RD algorithm.
  • Furthermore, we establish a quantitative relationship between ISAC signaling randomness and SAR imaging performance by adopting the normalized mean square error (NMSE) of a reference point target. Unlike conventional metrics that evaluate physical attributes in isolation, the NMSE provides a supplementary perspective that unifies the joint effects of signaling-induced sidelobes, peak energy loss, and noise amplification into a single indicator. This framework allows for a more direct characterization of how discrete QAM modulation impacts the overall reconstruction in data-aided SAR systems.
@@
Notation: Throughout the paper, $\mathbf{A}$, $\mathbf{a}$, and $a$ denote a matrix, vector, and scalar, respectively; $|a|$, $|\mathbf{A}|$ and $\frac{\mathbf{B}}{\mathbf{A}}$ represent the modulus of $a$, the element-wise modulus of $\mathbf{A}$ and the element-wise division of $\mathbf{B}$ by $\mathbf{A}$, respectively; $\mathbb{E}(\cdot)$, $tr(\cdot)$, $\|\cdot\|_{F}$, $(\cdot)^{T}$, $(\cdot)^{*}$, $(\cdot)^{H}$, $\odot$ and $\mathbf{I}$ denote the expectation, trace, Frobenius norm, transpose, conjugate, Hermitian, Hadamard (element-wise) product and identity matrix, respectively; A complex Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$ is represented as $\mathcal{CN}(\mu,\sigma^{2})$; Finally, $\mathrm{sinc}(x)=\frac{\sin(\pi x)}{\pi x}$.
Figure 1: Joint low-altitude OFDM-SAR imaging and data backhaul.
@@ II-A SAR Geometry
As depicted in Fig. 1, we consider a joint monostatic broadside stripmap UAV-SAR imaging embedded in continuous data backhaul scheme with cellular base stations in LAWNs. The UAV moves along the $y$-axis, with a height of $H_{p}$ and a constant velocity of $v$. To guarantee uninterrupted communications, the UAV’s coverage footprint must exceed the radius of a base station, ensuring at least one base station remains in line-of-sight of SAR beampattern at all times. Consequently, a quantitatively feasibility analysis is required.
Let us commence by analyzing the azimuth beamwidth and the elevation beamwidth, expressed as [5]
@@
where $\lambda$ is the wavelength, $D_{\mathrm{al}}$ and $D_{\mathrm{el}}$ represent azimuth and elevation antenna apertures, respectively. With a typical height $H_{p}$ from $300$ to $3000$ meters in LAWNs, the center of SAR elevation beamwidth defines the beam’s pointing elevation angle $\theta_{c}$. Accordingly, the azimuth coverage length $L_{\mathrm{az}}$ and the elevation coverage length $L_{\mathrm{el}}$ can be approximated as
@@
As clarified above, in order to guarantee the real-time data backhaul, $L_{\mathrm{az}}$ and $L_{\mathrm{el}}$ must be larger than the cellular coverage. For example, we consider $3.5$ GHz carrier frequency in sub-6 GHz band, azimuth and elevation antenna apertures as $D_{\mathrm{az}}=D_{\mathrm{el}}=0.1$ m and a central elevation angle $\theta_{c}=\frac{\pi}{4}$. At present, most commercial low-altitude UAV SAR platforms fly at around $H_{p}=1000$ m to balance spatial coverage and echoes’ magnitude. In this case, the corresponding ground coverage lengths are approximately $L_{\mathrm{az}}\approx 564$ m and $L_{\mathrm{el}}\approx 1899$ m, which obviously exceed the coverage of a 5G cellular cell. Based on these parameter analysis, the practical feasibility of integrating OFDM-SAR imaging into UAV data backhaul in LAWNs is geometrically convinced [^1] Low-altitude UAV can also employ a dual-beam mode: an imaging beam is utilized for ground sensing, while data backhaul is conducted through a dedicated directional beam pointing toward a fixed base station. This involves ISAC beam management issues [7], which are beyond the scope of this paper..
@@ II-B Transmit ISAC Signal Model
In (5), $s_{n,m}$ denotes the transmitted frequency-domain data, drawn from a finite alphabet such as 256-QAM constellation. In addition, $\Delta f=B_{r}/N=1/T_{p}$ represents the subcarrier interval in the frequency domain, and $\mathrm{rect}(t)$ represents the rectangle window, equal to $1$ for $0\leq t\leq 1$, and zero otherwise [^2] Note that the practical SAR has a sinc-like azimuth envelope [5, 45]. For the theoretical convenience, we use an ideal rectangle envelope instead.. Additionally, $T_{\mathrm{sym}}=T_{p}+T_{cp}$, where $T_{cp}$ denotes the length of CP. In order to eliminate the ISI and the inter-carrier interference (ICI) of received sensing signals, $T_{cp}$ must be larger than round-trip delay of the farthest target/path [9, 18, 19]. The total azimuth time is thus $T_{a}=MT_{\mathrm{sym}}$. Note that the power of a given codebook is normalized, i.e., $\mathbb{E}\{|s_{n,m}|^{2}\}=1$ for $\forall n,m$. This leads to the fact that the total average transmit energy of $M$ symbols is also normalized as $\mathbb{E}\{|x(t)|^{2}\}=\frac{1}{NM}\mathbb{E}\left\{\sum_{n,m}|s_{n,m}|^{2}\right\}=1$.